OpenStudy (blockcolder):

Prove that F(n)<2^n where F(n) is the nth Fibonacci number. Tried using strong induction, and failed. Am I doing something wrong?

OpenStudy (anonymous):

\(f(1) = 1 < 2^1\) and \(f(2) = 1 < 2^2 \). So it holds for \( n ≤ 2\) .

OpenStudy (blockcolder):

My attempt: Base cases: f(1)=1<2 f(2)=1<4 Inductive Step: Inductive hypothesis is: For all positive integers j less than or equal to k, f(j)<2^j. f(k+1)=f(k)+f(k-1) <2^k+2^(k-1) =3*2^(k-1) But since 4*2^(k-1)>3*2^(k-1), I can't proceed.

OpenStudy (anonymous):

Let it holds for all \( n = k \) Then \( f(k+1) = f(k) + f(k-1) < 2^k + 2^{(k -1)} < 2^k + 2^k =2^{(k+1)} \)

OpenStudy (anonymous):

Note: \(2^k \gt 2^{(k-1)}\).

OpenStudy (blockcolder):

Oh yeah. Forgot about that. Thanks!

OpenStudy (anonymous):

Glad to help :)